OpenStudy (marinedude42):

What is the slope of the line below? A. zero B. negative C. undefined D. positive

OpenStudy (marinedude42):

hold on, ill attach the graph if i can

OpenStudy (texaschic101):

whats the line ?

OpenStudy (anonymous):

He would post it

OpenStudy (texaschic101):

can't you just type it.....we do not need the graph

OpenStudy (marinedude42):

the line is going straight up and the coordinates are (3,5), and (3,2)

OpenStudy (texaschic101):

first you find the slope by using the slope formula slope(m) = (y2 - y1) / (x2 - x1) (3,5) x1 = 3 and y1 = 5 (3,2) x2 = 3 and y2 = 2 now we sub slope(m) = (2 - 5) / (3 - 3) slope(m) = -3/0 when there is a zero in the denominator, it is a vertical line with an undefined slope.

OpenStudy (texaschic101):

when the x values are the same...in this case, they are both 3, then the line is vertical and has an undefined slope. Now, if the y values would have been the same, it would have been a horizontal line with a 0 slope

OpenStudy (marinedude42):

thats what i just got when i done the coordinates, i got -3, and 0

OpenStudy (texaschic101):

the slope is undefined if it is a vertical line, which it is

OpenStudy (texaschic101):

Here are some acronyms that might help you... VUX -- vertical line, undefined slope, represented as x = a constant. HOY -- horizontal line, 0 slope, represented as y = a constant

OpenStudy (marinedude42):

thanks for your help, i dont understand some of this stuff

OpenStudy (texaschic101):

you see, with the two sets of points you were given...(3,5) and (3,2)...if you had to figure the line, it would be x = 3

OpenStudy (marinedude42):

what if the numbers were all negative, it would be -3 right?

OpenStudy (texaschic101):

if both x values were -3, then yes, the line would be x = -3 but the slope would still be undefined if there is a zero in the denominator.

OpenStudy (marinedude42):

oh, ok

OpenStudy (texaschic101):

just remember, vertical lines have a zero in the denominator and their slope is undefined. Horizontal lines have a zero in the numerator and their slope is 0

OpenStudy (marinedude42):

if its 7,0, then its undefined?

OpenStudy (texaschic101):

yep...because you got a zero in the denominator

OpenStudy (marinedude42):

ok, i get some of it now

OpenStudy (texaschic101):

wait...did you mean slope 7/0....if so, it is undefined.

OpenStudy (marinedude42):

yea, thats what i meant, sorry

OpenStudy (texaschic101):

no problem...do you have any questions

OpenStudy (marinedude42):

if i have the coordinates (-4,-6) and (2,-8) would it be horizontal,vertical,up and to the right, or down and to the right?

OpenStudy (marinedude42):

you are really good at math

OpenStudy (texaschic101):

slope(m) = (-8 - (-6) / (2 - (-4) slope(m) = (-8 + 6) / (2 + 4) slope(m) = -2/6 slope(m) = -1/3 DOWN AND TO THE RIGHT because the -1 is negative, it goes down, and 3 is positive, it goes to the right

OpenStudy (marinedude42):

youre cool, i can correst it later

OpenStudy (texaschic101):

if I had $1 for every mistake I made, I would be rich....thats so sad

OpenStudy (marinedude42):

its OK, i don't hold it against you

OpenStudy (texaschic101):

lol...good to know

OpenStudy (texaschic101):

do you have any more questions while I am still here ?

OpenStudy (marinedude42):

i dont hold a grudge or anything just cause you got the wrong answer, everybody makes mistakes

OpenStudy (texaschic101):

I am just glad I caught it

OpenStudy (marinedude42):

i medaled you anyway, so ur cool

OpenStudy (texaschic101):

thank you....if you have more problems, just tag me and I will see what I can do :)

OpenStudy (marinedude42):

thans

OpenStudy (marinedude42):

thanks

OpenStudy (texaschic101):

:) anytime

OpenStudy (marinedude42):

(5,0), (5,-9) up ant to the right or up and to the left?

OpenStudy (texaschic101):

look at your x values...they are both 5 when the x values are the same, the line is vertical with an undefined slope

OpenStudy (texaschic101):

if the y values were the same, the line would be horizontal with a 0 slope

OpenStudy (marinedude42):

i got another ? i dont mean to bother you though

OpenStudy (texaschic101):

what ya got ?

OpenStudy (marinedude42):

hold on a sec

OpenStudy (texaschic101):

no hurry..I am just sitting here with my bowl of cereal ....its really good too

OpenStudy (texaschic101):

just tag me

OpenStudy (marinedude42):

i have (-4,-6) and (2,-8)

OpenStudy (texaschic101):

what are you looking for ? The slope or the line ?

OpenStudy (marinedude42):

what direction is the line going?

OpenStudy (texaschic101):

slope = (y2 - y1) / (x2 - x1) (-4,-6) x1 = -4 and y1 = -6 (2,-8) x2 = 2 and y2 = -8 now we sub slope = (-8 - (-6) / (2 - (-4) slope = (-8 + 6) / (2 + 4) slope = -2/6 slope = -1/3 DOWN AND TO THE RIGHT

OpenStudy (marinedude42):

i thought so since most of the number were negative

OpenStudy (texaschic101):

In the slope, if the numerator is negative, the line goes down, if it is positive, the line goes up. In the denominator, if it is positive, it goes right, if it is negative, it goes left

OpenStudy (marinedude42):

you are really freaking smart

OpenStudy (texaschic101):

lol....not on some things.....not real good with geometry

OpenStudy (marinedude42):

thats what ive been asking you is geometry

OpenStudy (marinedude42):

solve 2|3(x-5)|+3=9

OpenStudy (texaschic101):

I just did that and sent it to her

OpenStudy (marinedude42):

ok, thanks

OpenStudy (texaschic101):

I just hope I did it right.....absolute values can be a pain in the butt

OpenStudy (marinedude42):

im not really smart in math, i owe you one.

OpenStudy (marinedude42):

In what direction is the line containing the points (-4, -6) and (2, -8) going? A. down and to the right B. vertical C. up and to the right D. horizontal